SPIN-ORBIT CONFIGURATION INTERACTION STUDY ON CUO

\underline{TIANXIAO YANG}, RAJNI TYAGI, RUSSELL M. PITZER and BRUCE E. BURSTEN, Department of Chemistry, The Ohio State University, 100 W.18th Avenue, Columbus, OH, 43210.

The ground and excited electronic states of linear CUO have been computed using relativistic spin-orbit configuration interaction as implemented in the COLUMBUS suite of programs. Four sets of geometries optimized by ADF^{a} and CASPT2^{b} method have been investigated. ADF and CASPT2 singlet geometries give that ^{1}Σ_{0}^{+} is 0.198 eV and 0.181 eV lower than ^{3}Φ_{2} respectively. ADF and CASPT2 triplet geometries give that ^{3}Φ_{2} is 0.306 eV and 0.455 eV lower than ^{1}Σ_{0}^{+} respectively. However, the lowest state at the triplet geometry is higher than the lowest state at the singlet geometry by 0.100 eV (ADF geometry) and 0.118 eV (CASPT2 geometry). Calculations^{b} by the CASPT2 method give the ^{3}Φ_{2} state lower in energy at all geometries. The experiments^{a} indicate that in the Ne matrix CUO is in the ^{1}Σ^{+} state and in the Ar matrix CUO is in the ^{3}Φ state. The electronic states of the system of CUO-noble gas atoms are also being investigated.

aJ. Li, B. E. Bursten, B. Liang, L. Andrews. Science \underline{295}, 2242, 2002.

bB. O. Roos, P. Widmark, L. Gagliardi. Faraday Discuss. \underline{124}, 57, 2003.